OpenStudy (anonymous):

FAN AND MEDAL The table below shows some values of f(x) and g(x) for different values of x. x f(x) = 2x + 24 g(x) = 5(0.5)x -2 20 -1 0 5 1 2.5 2 28 Complete the chart and determine the solution of the equation f(x) = g(x)? x = -2 x = -1 x = 1 x = 20

OpenStudy (ybarrap):

$$ f(x)=2x + 24 \\ g(x) = 5(0.5)x $$ For x=-1 $$ f(-1)=2(-1)+24=-2+24=22\\ g(-1)=5(0.5)(-1)=-2.5 $$ Similarly for x=0 $$ f(0)=?\\ g(0)=? $$ And so on... Now for g(x)=f(x) $$ 2x + 24=5(0.5)x\\ 2x-5(0.5)x=-24\\ x(2-2.5)=-24\\ x=-24/-1.5\\ =24/.5=48 $$ Does this help?

OpenStudy (anonymous):

I'm confused. What exactly does 24/.5 = 48 tell me ?

OpenStudy (ybarrap):

$$ x=24/.5=\cfrac{24}{0.5}=\cfrac{24}{\frac{1}{2}}=24\times2=48\\ $$ At this value of x $$ g(x)=f(x) $$ Because, $$ g(48)=5(0.5)(48)=120\\ f(48) = 2(48) + 24=120 $$ So $$ g(48)=f(48) $$

OpenStudy (anonymous):

So x=1 ?

OpenStudy (ybarrap):

no, x=48.

OpenStudy (anonymous):

f(-2) = 2(-2) + 24 f(-2) = 20 g(-2) = 5*4 g(-2) = 20 Its x = -2 because THEY BOTH EQUAL 20

OpenStudy (anonymous):

so its a or what